Can it be explained how you solve 16x-6(5/2 + 2)= 50? Im tried multiplying 5/2 by it's reciprocal. I'm just a little confused as to how to solve the problem.
Answer by stanbon(75887)

16x - 6(9/2)= 50
16x - 27 =50
16x = 77
x = 77/16.
Cheers,
Stan H.